A certain rifle bullet has a mass of 7.89 g. Calculate the de Broglie wavelength of the bullet traveling at 1313 miles per hour. Physical constants can be found here. Number λ = 4.929 × 10-29
Answer #1

λ = h/ P

P = m*v

here

m = 7.89*10^-3 kg

v= 1313 mph =586.96352 m / s so

P = m*v = 7.89*10^-3*586.96352 =4.63114217

λ = h/ P = 6.626*10^-34/ 4.63114217 =  1.43074856e-34 m answer